# Femtoscopy of $DN$ and $\bar{D}N$ systems

### v1

@@ Abstract
The capability of the ALICE@LHC and STAR@RHIC experiments to reconstruct $D$ mesons has enabled femtoscopic correlation measurements of open-charm mesons in both small and large systems. In this work, we present a theoretical calculation of the correlation functions of $D$ and $\bar{D}$ mesons with nucleons, based on the Koonin-Pratt formalism applied to the systems produced in these high-energy collisions. We use an effective Lagrangian that models the interaction between charmed mesons and baryons. From the unitarized on-shell amplitudes, we extract the low-energy scattering parameters and compute the correlation functions using the Lednický-Lyuboshitz approximation. Additionally, we apply the TROY formalism to obtain the off-shell $T$-matrix in coupled channels and reconstruct the full wave function, which we incorporate into the Koonin-Pratt formula. We compare these two approaches and provide predictions for different correlated pairs, including the effects of the Coulomb interaction and thermal weights for the coupled channels. These results can be tested against current and future experimental data from ALICE and STAR collaborations in both proton-proton and heavy-ion collisions.
@@ I Introduction
Over the past years, femtoscopy has turned out to be a promising tool to study the interactions among hadrons (see the review of Ref. [1]). Femtoscopy is based on determining hadron-hadron correlations in momentum space as the ratio between the distribution of relative momenta for pairs produced in the same collision and in different collisions. Since the pioneer works in [2, 3] and further developments of the STAR Collaboration [4, 5], the ALICE Collaboration has taken the leading role and determined the correlations for different hadron-hadron systems (see, for example, the works in [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]).
In the past years the ALICE Collaboration has started investigating hadron-hadron interactions in the charm sector using femtoscopic techniques. More precisely, the correlation functions of $D^{(\pm)}\pi^{(\pm)}$ and $D^{(\pm)}K^{(\pm)}$ [26] as well as $D^{-}p$ and $D^{+}\bar{p}$ [27] in high-multiplicity pp collisions at 13 TeV have been recently obtained, thus giving new insights into interactions involving with $D$ mesons.
@@ II.1 Meson-baryon strong interaction
An effective description of meson-baryon scattering involving charm and anticharm degrees of freedom is developed in Ref. [41]. Using that approach, we focus on the $C=1$ and $C=-1$ sectors, to accommodate the $DN$ and $\bar{D}N$ channels, respectively. We follow the notation and conventions given in Ref. [42].
@@
$N=\sqrt{\dfrac{M_{i}+E_{i}(\bm{p}_{i})}{2M_{i}}}\ \sqrt{\dfrac{M_{j}+E_{j}(\bm{p}_{j})}{2M_{j}}}\ ,$ (3)
is a normalization factor arising from the Dirac spinors; and $M_{i}$, $M_{j}$ and $E_{i}$, $E_{j}$ are the masses and energies of the baryons in channels $i$ and $j$, respectively. Equation (2) serves as an off-shell kernel for the meson-baryon scattering in the ZR approximation. The on-shell version of [42] is simply achieved by setting $p_{i/j}^{2}=E_{i/j}^{2}-M_{i/j}^{2}$. In this work, both off-shell and on-shell versions of this potential in the ZR approximation will be used.
@@
where the new isospin coefficients
@@
with a correction factor $\kappa=1/4$ only when the exchanged vector meson contains a charm quark, to balance from the universal vector mass assumption [42].
In this work, we focus on four isospin/strangess/charm sectors. To account for $D^{0}p$ and $D^{+}p$ cases we will consider the sector $\{ISC\}=\{0,0,1\}$ with 7 coupled channels,
@@
leaving aside heavier channels with additional $c\bar{c}$ content, and also the sector $\{ISC\}=\{1,0,1\}$ with 8 coupled channels,
@@
For the $D^{-}p$ and $\bar{D}^{0}p$ we work with the $\{ISC\}=\{0,0,-1\}$ and $\{ISC\}=\{1,0,-1\}$ sectors, each of them with a single channel, $\bar{D}N$.
@@ II.2 Off- and on-shell $T$-matrices
To regularize the ultraviolet (UV) limit of the perturbative potential in Eq. (8) we apply, as in Ref. [35], an exponential form factor,
$F(q^{\prime},q)=\exp\left(-\frac{q^{\prime 2}+q^{2}}{\Lambda^{2}}\right),$ (9)
to the potential, with the regulator $\Lambda$, universal for all scattering channels, to be fixed as we will explain later.
@@
In the on-shell version, the loop function $G_{l}(\sqrt{s})$ also needs to be regularized, which we will do by setting a UV 3-momentum cut-off, $k_{\textrm{max}}$ and using the analytical expression given in Ref. [53]. We will also use a universal value of $k_{\textrm{max}}$ for all $DN$ and $\bar{D}N$ channels.
@@ II.3 Femtoscopy correlation function
II.3.1 LednickĂ˝-Lyuboshitz approximation
@@
$C^{\textrm{C}}_{\textrm{LL}}(k)=A_{\textrm{C}}(k)\left\{1+\dfrac{|f_{\textrm{C}}(q)|^{2}}{4r_{0}^{2}}\left[1+e^{-(2kr_{0})^{2}}+A_{\textrm{C}}^{2}(1-e^{-(2kr_{0})^{2}})-\dfrac{d_{0}}{\sqrt{\pi}r_{0}}\right]\dfrac{2\ {\rm Re}f_{\textrm{C}}(k)}{\sqrt{\pi}r_{0}}F_{1}(2kr_{0})-\dfrac{A_{\textrm{C}}\ {\rm Im}f_{\textrm{C}}(k)}{r_{0}}F_{2}(2kr_{0})\right\}\ ,$ (14)
@@
where $\gamma=Z_{1}Z_{2}\mu\alpha/k$ is the Sommerfeld factor, and $a_{0}$ and $d_{0}$, are the scattering length and effective range, respectively. The scattering length and effective range will be obtained from the on-shell $T$ matrix. In our convention, $a_{0}>0$ indicates attraction, while $a_{0}<0$ indicates repulsion or the presence of a bound state. Finally, the function $h(x)$ is defined as,
@@
where $\psi(z)$ is the digamma function. In the absence of Coulomb interaction, $\gamma\rightarrow 0$, we recover the standard LL correlation function [58],
@@ II.3.2 Off-shell $T$-matrix (TROY) approach
where $m_{\perp}=\sqrt{m^{2}+p_{\perp}^{2}}$, $m$ is the mass of the particle, $p_{\perp}$ its transverse momentum, $y$ is the rapidity, $\varphi$ is the azimuthal angle of the particle’s momentum, and $T$ is the temperature. We integrate this multiplicity without kinematic restrictions and use $T=171$ MeV for high-multiplicity pp collisions at $\sqrt{s}=13$ TeV at the LHC [22]. Then, the thermal weight $w_{i}$ for a given pair—using $w_{f}=1$ as the reference value—is given by
@@
$\varphi_{if}(k;r)$ $=j_{0}(kr)\delta_{if}$ (24)
+∫$+\int\frac{d^{3}q^{\prime}}{(2\pi)^{3}}\frac{2M_{i}(E_{i}+\omega_{i})}{2E_{i}\omega_{i}}\frac{T_{if}(q^{\prime},k;\sqrt{s})j_{0}(q^{\prime}r)}{s-(E_{i}+\omega_{i})^{2}+i\eta}\ ,$
where $j_{0}(x)$ is the spherical Bessel function of zeroth order, and $\sqrt{s}=\sqrt{M_{f}^{2}+k^{2}}+\sqrt{m_{f}^{2}+k^{2}}$.
@@ III Results
In this section, we present the correlation function of $DN$ and $\bar{D}N$ pairs, obtained using different approximations. In Sec. III.1 we use the LL model with Coulomb interactions for small systems. In Sec. III.2 we apply the off-shell TROY formalism for the same small systems. In Sec. III.4 we use TROY to compare to ALICE results in pp collisions and predict the result for larger systems, as those studied by the STAR collaboration.
III.1 LL approximation
For the LL case, we apply the on-shell results by solving (10) using first the ZR approximation with $g=6.0$ and $k_{\textrm{max}}=674$ MeV. Then, for the WT interaction we use $f=1.15f_{\pi}$ with $f_{\pi}=92$ MeV and $k_{\textrm{max}}=752$ MeV.
@@ III.1 LL approximation
First, we start by presenting the scattering lengths in the isospin basis for our model, while comparing our results with the outcome from other works. The scattering lengths are displayed in Tables 1 and 2 for $DN$ and $\bar{D}N$, respectively.
@@
With regard to the $DN$ system, the $s$-wave $I=0$ and $I=1$ scattering lengths predicted by the WT model and those in the ZR approximation, both in the on-shell case, are very similar. This is an expected result, as the difference between the two approaches arises because of the slightly different energy-dependent $DN$ bare interaction. In both WT and ZR cases, the $I=0$ $\Lambda_{c}(2595)$ is generated dynamically at its mass by adjusting the cutoff, whereas in the $I=1$ sector a $\Sigma_{c}$ with a mass close to the $DN$ threshold of 2806 MeV is obtained, thus leading to a large scattering length. Note that the scattering lengths reported in Ref. [46] for the WT model of Ref. [44] slightly differ from those of the WT model shown here due to the slightly different cutoff used to regularize the loop function. Moreover, we show the scattering lengths of the $DN$ interaction of Ref. [41], reported in Ref. [43].
In that work, there is no $I=1$ quasi-bound state close to the $DN$ threshold but lies 180 MeV below, so the size of the corresponding scattering length is smaller than those from the WT and ZR approaches. As for the scattering lengths coming from the meson-exchange model of Ref. [46], they are qualitatively similar in both isospin sectors to those shown for WT and ZR models. This is due to the fact that the meson-exchange model also generates dynamically the $I=0$ $\Lambda_{c}(2595)$ and the $I=1$ $\Sigma_{c}(2800)$ for masses close to those obtained in our present study. And finally, for completeness, we show the scattering lengths for the SU(8) model of Ref. [45], reported in Ref. [46]. In that approach, both $I=0$ and $I=1$ scattering lengths are positive, being the $I=0$ one compatible with zero. The discrepancy with previous works has its origin in the different and richer resonant-hole composition close to $DN$ threshold, as described in Ref. [45].
@@
In the case of $\bar{D}N$, the $I=0$ scattering length in the WT approach is zero, as the corresponding interaction is strictly vanishing in this channel. A result compatible with zero is consistently obtained in the ZR approach, as seen in Table 2. For $I=1$, both the WT and ZR prescriptions produce a similarly sizable negative scattering length, thus indicating a repulsive $\bar{D}N$ interaction in that sector. Upon comparing these results with those from Ref. [43], we observe that, while the $I=1$ value is very close to the ones for WT and ZR, there is a disagreement in the $I=0$ scattering length, that we are not able to trace back to the regularization method. In Ref. [47] a value close to zero was obtained for the $I=0$ scattering length, whereas a more repulsive scattering length than our results for WT and ZR was determined for $I=1$. This model is based on the exchange of mesons between $\bar{D}$ and $N$ supplemented by a short-range one-gluon exchange and, in fact, about half of the repulsive scattering length in $I=1$ comes from the hadronic meson-exchange contributions, which can be mapped to a certain extent to the WT interaction used in the present work. This result was revisited in Ref. [27], by modifying the coupling to the scalar meson, leading to a much more attractive interaction in $I=0$ and a value close to zero for $I=1$. Furthermore, the analysis of Ref. [49], with a model that considers a long range one-pion exchange interaction, obtains a bound $\bar{D}N$ state in $I=0$, leading to a very large and negative $I=0$ scattering length, whereas the $I=1$ one is much smaller. Notice that the value of Ref. [49] for the $I=0$ scattering length is quoted in [27] with an opposite sign, because the convention in the former paper is contrary to the ALICE one (and ours). The work of Ref. [48] uses a quark model that confines color and realizes dynamical chiral symmetry breaking, giving rise to values for $I=0$ and $I=1$ closer to the WT and ZR results, as reported in Ref. [27].
$DN$ $a_{0}~{}({I=0})[{\rm fm}]$ $a_{0}~{}(I=1)[{\rm fm}]$
[…]
Hofmann-Lutz [41, 43] −$-0.43$$-0.41$
[…]
@@
Table 1: Scattering lengths for $DN$ with the Weinberg-Tomozawa (WT) and Zero-Range (ZR) approximations compared with other theoretical models in the literature.
@@
WT 00 −$-0.29$
ZR $0.02$$-0.34$
Hofmann-Lutz [43] −$-0.16$$-0.26$
Haidenbauer [47] $0.07$$-0.45$
@@
We now present the correlation functions for charged particles, obtained using the scattering lengths from this work and compare them with those obtained using the scattering lengths from previous studies, as discussed earlier. In all cases, we have used the LL approximation with Coulomb interaction of Eq. (14), employing a source radius of $r_{0}=1~{}\rm{fm}$, and a vanishing scattering range. We have checked in the $\bar{D}N$ case that the effect of the scattering range in the LL formula is negligible.
The correlation functions for the purely $I=1$ $D^{+}p$ system are shown in Fig. 1. The repulsive effect caused by the Coulomb interaction, leading to a correlation function below one, dominates the correlation function, although some differences can be appreciated between the models. The negative sign of the scattering length simulates a repulsive interaction in the majority of models, reducing further the correlation function with respect the the Coulomb only contribution. As mentioned before, the model of Ref. [45] presents a positive scattering length, hence an attractive interaction, leading to a higher correlation function compared to the other models and reaching values larger than one for a certain range of momenta.
The $D^{-}p$ correlation functions for different models are shown in Fig. 2. In this case, the attractive Coulomb effect dominates, leading in general to values of the correlation function larger than one. Since this system consists of a combination of isospins, the scattering length used is taken as $a=(a_{I=0}+a_{I=1})/2$ for each model. The variation in the scattering lengths presented earlier results in significantly different correlation functions. The small and negative averaged scattering lengths obtained in this work with the WT and ZR models lead to correlation functions that are very similar to those from other studies. The repulsive interaction associated with these values causes only a small deviation compared to the dominant attractive Coulomb effect. In contrast, the large negative scattering length reported in Ref. [49] for $I=0$ induces a substantial apparently repulsive effect, as seen in the dotted blue line in Fig. 2, where the correlation function is significantly lower than in the other works. This behavior is in fact signalling the presence of a bound state generated by the model of Ref. [49]. The opposite effect is observed in the results revisited in Ref. [27] using the model from Ref. [47]. In this case, the positive and large scattering length indicates an attractive interaction, leading to larger correlation function that agrees with experimental data [27].

### v2

@@ Abstract
The capability of the ALICE@LHC and STAR@RHIC experiments to reconstruct $D$ mesons has enabled femtoscopic correlation measurements of open-charm mesons in both small and large systems. In this work, we present a theoretical calculation of the correlation functions of $D$ and $\bar{D}$ mesons with nucleons, based on the Koonin-Pratt formalism. We employ an effective Lagrangian to model the interaction between charmed mesons and baryons and apply the TROY formalism to obtain the off-shell $T$-matrix in coupled channels, incorporating the effect of the Coulomb interaction when the pair involves two charged particles. The resulting full coupled-channel wave function is inserted into the Koonin-Pratt equation with channel weights derived from a thermal model. Additionally, we compute the correlation functions using the Lednický-Lyuboshitz approximation with low-energy scattering parameters extracted from the unitarized amplitudes. We compare these two approaches and provide predictions for different correlated pairs. Our results can be tested against current and future experimental data from the ALICE and STAR collaborations in both proton-proton and heavy-ion collisions.
@@ I Introduction
Over the past years, femtoscopy has turned out to be a promising tool to study the interactions among hadrons (see the review of Ref. [1]). Femtoscopy is based on determining hadron-hadron correlations in momentum space as the ratio between the distribution of relative momenta for pairs produced in the same collision and in different collisions. Since the pioneer works in [2, 3] and further developments of the STAR Collaboration [4, 5, 6], the ALICE Collaboration has taken the leading role and determined the correlations for different hadron-hadron systems (see, for example, the works in [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]).
In the past years the ALICE Collaboration has started investigating hadron-hadron interactions in the charm sector using femtoscopic techniques. More precisely, the correlation functions of $D^{(\pm)}\pi^{(\pm)}$ and $D^{(\pm)}K^{(\pm)}$ [27] as well as $D^{-}p$ and $D^{+}\bar{p}$ [28] in high-multiplicity pp collisions at 13 TeV have been recently obtained, thus giving new insights into interactions involving $D$ mesons.
@@ II.1 Meson-baryon strong interaction
An effective description of meson-baryon scattering involving charm and anticharm degrees of freedom is developed in Ref. [42]. Using that approach, we focus on the $C=1$ and C=1C=-1 sectors, to accommodate the $DN$ and $\bar{D}N$ channels, respectively. We follow the notation and conventions given in Ref. [43].
@@
$N=\sqrt{\dfrac{M_{i}+E_{i}}{2M_{i}}}\ \sqrt{\dfrac{M_{j}+E_{j}}{2M_{j}}}\ ,$ (3)
is a normalization factor arising from the Dirac spinors; and $M_{i}$, $M_{j}$ and $E_{i}=(s+M_{i}^{2}-m_{i}^{2})/(2\sqrt{s})$, $E_{j}=(s+M_{j}^{2}-m_{j}^{2})/(2\sqrt{s})$ are the masses and energies of the baryons in channels $i$ and $j$, respectively. Equation (2) serves as an off-shell kernel for the meson-baryon scattering in the ZR approximation. The on-shell version of [43] is simply achieved by setting $p_{i/j}^{2}=E_{i/j}^{2}-M_{i/j}^{2}$. In this work, both off-shell and on-shell versions of this potential in the ZR approximation will be used.
@@
where the new isospin coefficients,
@@
carry a correction factor $\kappa=1/4$ only when the exchanged vector meson contains a charm quark, to balance from the universal vector mass assumption [43].
In this work, we focus on four isospin/strangess/charm sectors. To account for $D^{0}p$ and $D^{+}p$ cases we will consider the $\{ISC\}=\{0,0,1\}$ DNDN sector with 7 coupled channels,
@@
leaving aside heavier channels with additional $c\bar{c}$ content, and also the $\{ISC\}=\{1,0,1\}$ DNDN sector with 8 coupled channels,
@@
For the $D^{-}p$ and $\bar{D}^{0}p$ cases we work with the $\{ISC\}=\{0,0,-1\}$ and $\{ISC\}=\{1,0,-1\}$ $\bar{D}N$ sectors, each of them having a single channel.
@@ II.2 Off- and on-shell $T$-matrices
Both this loop integral and that of Eq. (8) exhibit an ultraviolet (UV) divergence, which we will regularize by introducing a UV 3-momentum cutoff, $k_{\textrm{max}}$, whose value will be determined by imposing the position of the dynamically generated $\Lambda_{c}(1295)$ in the $DN$ sector. The same value will be used in the $\bar{D}N$ sector
@@ II.3 Femtoscopy correlation function
II.3.1 Lednický-Lyuboshitz approximation
@@
$C^{\textrm{C}}_{\textrm{LL}}(k)=A_{\textrm{C}}(k)\left\{1+\dfrac{|f_{\textrm{C}}(q)|^{2}}{4r_{0}^{2}}\left[1+e^{-(2kr_{0})^{2}}+A_{\textrm{C}}^{2}(1-e^{-(2kr_{0})^{2}})-\dfrac{d_{0}}{\sqrt{\pi}r_{0}}\right]+\dfrac{2\ {\rm Re}f_{\textrm{C}}(k)}{\sqrt{\pi}r_{0}}F_{1}(2kr_{0})-\dfrac{A_{\textrm{C}}\ {\rm Im}f_{\textrm{C}}(k)}{r_{0}}F_{2}(2kr_{0})\right\}\ ,$ (13)
@@
where $\gamma=Z_{1}Z_{2}\mu\alpha/k$ is the Sommerfeld factor. The parameters $a_{0}$ and $d_{0}$, stand for the scattering length and effective range, respectively, and determine the low-energy expansion of the strong-only scattering amplitude. In our convention, $a_{0}>0$ indicates attraction, while $a_{0}<0$ indicates repulsion or the presence of a bound state. Finally, the function $h(x)$ is defined as,
@@
where $\psi(z)$ is the digamma function. In the absence of Coulomb interaction, on sets $\gamma\rightarrow 0$ and recovers the standard LL correlation function [58],
@@ II.3.2 Off-shell $T$-matrix (TROY) approach
where $_{\perp}=\sqrt{m^{2}+p_{\perp}^{2}}, mm$ is the mass of the particle, $p_{\perp}$ its transverse momentum, $y$ is the rapidity, $\varphi$ is the azimuthal angle of the particle’s momentum, and $T$ is the temperature. We integrate this multiplicity without kinematic restrictions and use $T=171$ MeV for high-multiplicity pp collisions at $\sqrt{s}=13$ TeV at the LHC [23]. Then, the thermal weight $w_{i}$ for a given pair—using $w_{f}=1$ as the reference value—is given by
@@
$\varphi_{if}(k;r)$ $=$ $j_{0}(kr)\delta_{if}$
$+$ $\int\frac{d^{3}q^{\prime}}{(2\pi)^{3}}\frac{2M_{i}(E_{i}+\omega_{i})}{2E_{i}\omega_{i}}\frac{T_{if}(q^{\prime},k;\sqrt{s})j_{0}(q^{\prime}r)}{s-(E_{i}+\omega_{i})^{2}+i\eta}\ ,$
where $j_{0}(x)$ is the spherical Bessel function of zeroth order, and $\sqrt{s}=\sqrt{M_{f}^{2}+k^{2}}+\sqrt{m_{f}^{2}+k^{2}}$. We note that, in the case of two charged particles, the $T$ matrix entering the above equation is solved including both the strong and Coulomb interaction kernels, as shown in Ref. [36] and described in Appendix A.
For comparison, we also compute the correlation function given in Eq. (19) using the on-shell approach, where the interacting wave function of Eq. (II.3.2) is obtained from the on-shell form of the $T$ matrix.
@@ III Results
In this section, we present the correlation function of $DN$ and $\bar{D}N$ pairs, calculated using different approximations. In Sec. III.1 we employ the LL approach for small system sizes and compare our results with those obtained for other interaction models in the literature. In Sec. III.2 we apply the off-shell TROY formalism for the same small systems. In Sec. III.4 we use TROY to compare to ALICE results in pp collisions and predict the correlation functions for larger systems, as those studied by the STAR collaboration.
III.1 LL calculations
In this section we present results for the correlation functions of charged particle pairs, $D^{+}p$ and $D^{-}p$, within the LL approach of Eq. (13). We derive the required low-energy parameters from our on-shell version of the $T$-matrix given in Eq. (9). For the ZR interaction kernel we employ $g=6.0$ and $k_{\textrm{max}}=674$ MeV, while we use $f=1.15f_{\pi}$ with $f_{\pi}=92$ MeV and $k_{\textrm{max}}=752$ MeV in the case of the WT interaction.
@@ III.1 LL calculations
First, we present the $DN$ and $\bar{D}N$ scattering lengths in the isospin basis in Tables 1 and 2, respectively, which also display the outcome from other works for comparison.
@@
With regard to the $DN$ system, the $s$-wave $I=0$ and $I=1$ scattering lengths predicted by the WT model and those in the ZR approximation, both in the on-shell case, are very similar. This is an expected result, as the difference between the two approaches arises because of the slightly different energy-dependent $DN$ interaction kernel. In both WT and ZR cases, the $I=0$ $\Lambda_{c}(2595)$ is generated dynamically at its mass by adjusting the cutoff, whereas in the $I=1$ sector a $\Sigma_{c}$ with a mass close to the $DN$ threshold of 2806 MeV is obtained, thus leading to a scattering length having a large negative real part. Note that the scattering lengths reported in Ref. [47] for the WT model of Ref. [45] slightly differ from those of the WT model shown here due to the slightly different cutoff used to regularize the loop function. Moreover, we show the scattering lengths of the $DN$ interaction of Ref. [42], reported in Ref. [44]. In that work, there is no $I=1$ quasi-bound state close to the $DN$ threshold but it lies 180 MeV below, explaining the smaller size of the corresponding scattering length compared to the WT and ZR results shown here. As for the scattering lengths coming from the meson-exchange model of Ref. [47], they are qualitatively similar in both isospin sectors to those shown for WT and ZR models. This is due to the fact that the meson-exchange model also generates dynamically the $I=0$ $\Lambda_{c}(2595)$ and the $I=1$ $\Sigma_{c}(2800)$ for masses close to those obtained in our present study. We also show the scattering lengths for the SU(8) model of Ref. [46], reported in Ref. [47]. In that approach, both $I=0$ and $I=1$ scattering lengths are positive, being the $I=0$ one compatible with zero. The discrepancy with previous works has its origin in the presence of other resonant states beyond the $I=0$ $\Lambda_{c}(2595)$ and the $I=1$ $\Sigma_{c}(2800)$ close to the $DN$ threshold, as described in Ref. [46]. Finally, we include the results of [62] where the scattering lengths were extracted from experimental data of the $pD^{0}$ invariant-mass distribution in the three-body decay $\Lambda_{b}\rightarrow\pi^{-}pD^{0}$. While the resulting $I=0$ scattering length covers roughly the range of values obtained by the various theoretical models, the central value of the $I=1$ scattering length turns out to have more sizable real and imaginary parts. More notably, even taking the uncertainties of the analysis of Ref. [62] into account, the real part of the $I=1$ scattering length cannot be reproduced by the models.
@@
In the case of $\bar{D}N$, the $I=0$ scattering length in the WT approach is zero, as the corresponding interaction is strictly vanishing in this channel. A result compatible with zero is consistently obtained in the ZR approach, as seen in Table 2. For $I=1$, both the WT and ZR prescriptions produce a similarly sizable negative scattering length, thus indicating a repulsive $\bar{D}N$ interaction in that sector. Upon comparing these results with those from Ref. [44], we observe that, while the $I=1$ value is very close to the ones for WT and ZR, there is a disagreement in the $I=0$ scattering length, that we are not able to trace back to the regularization method. In Ref. [48] a value close to zero was obtained for the $I=0$ scattering length, whereas a more repulsive scattering length than our results for WT and ZR was determined for $I=1$. This model is based on the exchange of mesons between $\bar{D}$ and $N$ supplemented by a short-range one-gluon exchange and, in fact, about half of the repulsive scattering length in $I=1$ comes from the hadronic meson-exchange contributions, which can be mapped to a certain extent to the WT interaction used in the present work. This result was revisited in Ref. [28], by modifying the coupling to the scalar meson, leading to a much more attractive interaction in $I=0$ and an almost negligible one in $I=1$. Furthermore, the analysis of Ref. [50], with a model that considers a long range one-pion exchange interaction, obtains a bound $\bar{D}N$ state in $I=0$, leading to a very large and negative $I=0$ scattering length, whereas the $I=1$ one is much smaller. Notice that the value of Ref. [50] for the $I=0$ scattering length is quoted in [28] with an opposite sign, because the convention in the former paper is contrary to the ALICE one (and ours). The work of Ref. [49] uses a quark model that confines color and realizes dynamical chiral symmetry breaking, giving rise to values for $I=0$ and $I=1$ closer to the WT and ZR results, as reported in Ref. [28].
DNDN $a_{0}~({I=0})[{\rm fm}]$ $a_{0}~(I=1)[{\rm fm}]$
WT 0.58-0.58 $1.55+\rm{i}0.75$
ZR 0.63-0.63 $1.79+\rm{i}0.32$
Mizutani [45, 47] 0.57-0.57 $1.47+\rm{i}0.65$
Hofmann-Lutz [42, 44] 0.43-0.43 0.41-0.41
Haidenbauer [47] $0.41+\rm{i}0.04$ $2.07+\rm{i}0.57$
Garcia-Recio [46, 47] $(4+\rm{i}2)\cdot 10^{-3}$ $0.33+\rm{i}0.05$
Sakai [62] $0.79_{-0.61}^{+0.66}$ $3.8_{-2.0}^{+1.4}+\rm{i}2.7_{-2.7}^{+1.6}$
@@
Table 1: Scattering lengths for DNDN with the Weinberg-Tomozawa (WT) and Zero-Range (ZR) approximations compared with other theoretical models in the literature.
@@
WT 0 0.29-0.29
ZR $0.02$ 0.34-0.34
Hofmann-Lutz [44] 0.16-0.16 0.26-0.26
Haidenbauer [48] $0.07$ 0.45-0.45
@@
We now present the correlation functions for charged particle pairs, using the LL approximation with Coulomb interaction of Eq. (13) and employing a source radius of $r_{0}=1~\rm{fm}$. Since the aim is to present a comparison of our results with those obtained for the interaction models of previous studies, which only provide information on the scattering length, our results are obtained setting the scattering range to zero. We have checked that, in the $\bar{D}N$ case, the effect of the scattering range in the LL formula is negligible.
The correlation functions for the purely $I=1$ $D^{+}p$ system are shown in Fig. 1. The repulsive effect caused by the Coulomb interaction, leading to a correlation function below one, dominates the correlation function, although some differences can be appreciated between the models. The negative sign of the scattering length simulates a repulsive interaction in the majority of models, reducing further the correlation function with respect the the Coulomb only contribution. As mentioned before, the model of Ref. [46] presents a positive scattering length, hence an attractive interaction, leading to a higher correlation function compared to the other models and reaching values larger than one for a certain range of momenta. The scattering lengths in [62], in particular the large value of the $I=1$ channel, produce a correlation function peaked around $k=30$ MeV and above one, signalling the strong coupling of two near-threshold bound states found in that reference, identified to a possible two-pole structure of the $\Sigma_{c}(2800)$ state.
The $D^{-}p$ correlation functions for different models are shown in Fig. 2. In this case, the attractive Coulomb effect dominates, leading in general to values of the correlation function larger than one. Since this system consists of a combination of isospins, the scattering length used is taken as $a=(a_{I=0}+a_{I=1})/2$ for each model. The dispersion of scattering length values presented earlier results in significantly different correlation functions. The small and negative averaged scattering lengths obtained in this work with the WT and ZR models lead to correlation functions that are very similar to those from other studies. The repulsive interaction associated with these values causes a moderate deviation compared to the dominant attractive Coulomb effect. In contrast, the large negative scattering length reported in Ref. [50] for $I=0$ induces a substantial apparently repulsive effect, as seen in the dotted blue line in Fig. 2, where the correlation function is significantly lower than in the other works. This behavior is in fact signalling the presence of a bound state generated by the model of Ref. [50]. The opposite effect is observed in the results revisited in Ref. [28] using the model from Ref. [48]. In this case, the positive and large scattering length indicates an attractive interaction, leading to larger correlation function that agrees with experimental data [28].